# FastSLM: Hierarchical Frame Q-Former for Effective Speech Modality Adaptation

### v1

@@ 4.1 Dataset Description
Pre-training Dataset: As described in Section 3.2, we constructed a bilingual dataset comprising 15K speech-text pairs to adapt ASR capabilities to the LLM during the pre-training stage. Including 9,152 hours of English speech-text pair (LibriSpeech (Panayotov et al., 2015), GigaSpeech-L (Chen et al., 2021), Voxpopuli (Wang et al., 2021), SpgiSpeech-M (O’Neill et al., 2021), Earnings-22 (Rio et al., 2022), AMI (Kraaij et al., 2005), Common Voice 15 (Ardila et al., 2019), AI-HUB ASR-En (The Open AI Dataset Project, 2021)) dataset, and 7,812 hours of Korean speech-text pair (AI-Hub-ASR-Ko (The Open AI Dataset Project, 2021)) dataset. A detailed description of the pre-training dataset can be found in Appendix D Table 7.
Long-form speech Dataset: To enhance the model capacity to process long-form speech input, particularly for tasks such as SSUM and SQQA, we constructed a dedicated long-form speech dataset. For English, we curated a total of 219 hours of long-form speech. For Korean, we curated a total of 384 hours of long-form speech.
@@
  • AST: We evaluated En2Ko and Ko2En translation on the Fleurs, and Minds14 (Gerz et al., 2021) dataset. In addition to the standard BLEU score (Post, 2018), we employed a GPT-4 based evaluation to overcome BLEU limitations with semantically valid but lexically different translations (Zhao et al., 2023b). Please refer to Appendix E for the AST judge prompt for GPT-4.
  • SSUM: Evaluation was conducted on SDS-PART6 and KMSS-speech. Summarization quality was assessed using GPT-4 scoring with the LLM-as-a-judge framework (Zheng et al., 2023). Please refer to Appendix E for the SSUM judge prompt for GPT-4.
@@ 4.2 Experimental Setup
Model Architecture: FastSLM employs the encoder from Whisper-large-v3 (Radford et al., 2023) for speech feature extraction and adopts Qwen3-4B (Yang et al., 2025) as the backbone LLM for text generation. Despite its relatively compact size, Qwen3-4B exhibits sufficient capacity for comprehending speech-derived representations. In contrast to prior SLMs that typically utilize LLM backbone with 7 to 14 billion (B) parameters (Chu et al., 2023; 2024; Tang et al., 2024; Yu et al., 2024; Rubenstein et al., 2023; Ding et al., 2025; Liu et al., 2025a), FastSLM achieves a favorable cost-performance trade-off by leveraging lightweight architecture without compromising performance (Abouelenin et al., 2025; Ghosh et al., 2025). The HFQ-Former module within FastSLM compresses frame-level features via a hierarchical query-based mechanism. The number of learnable queries of $\mathbf{Q}_{i}^{A}$ was set to 80 cost-effectively (Yu et al., 2024), and the number of learnable queries of $\mathbf{Q}_{c}^{A}$ used as contextual input in LLM was set to 50 through the experiment. For further ablation studies and design justifications, please refer to Section 4 Fig. 3.
@@
The model implementation details and the training setup are summarized in Appendix G.
@@ 4.3 Comparison with Baseline Q-Former
HFQ-Former (ours) $2.09$ $4.67$ $\mathbf{6.99}$ $\mathbf{1.67}$ $\mathbf{2.51}$
@@
In contrast, HFQ-Former achieves a strong balance between accuracy and efficiency: it reduces the token rate by 37% compared to WQ-Former and lowers LLM FLOPs by 31.2% (3.65T → 2.51T), while still improving WER on VoxPopuli. These results indicate that HFQ-Former provides a significantly more efficient speech-to-LLM alignment mechanism, greatly reducing the computational burden of autoregressive decoding for long-form speech without compromising recognition quality.
@@ 4.4 Quantitative Results
Table 3: Comparison of FastSLM with other SLMs on various tasks. WER is lower than best, accuracy (ACC), BLEU, and score higher than best. N/A indicates the model does not have such a capability. ‘*’ indicates results fine-tuned on an additional Korean dataset provided in the official Hugging Face supplementary material. Detailed ASR benchmark WER for each dataset is reported in Appendix H.
#speech tokens/30 Sec. 50 1500 101 375 750 960
ASR (En) WER $\downarrow$ OpenASR $6.83$ $7.44$ $7.43$ $\mathbf{6.14}$ $7.05$ $9.29$
ASR (Ko) CER $\downarrow$ Fleurs Common Voice 15 $\mathbf{3.82}$ $7.92$ N/A N/A N/A $4.55$
AST (En2Ko) BLEU/Score (1-5) $\uparrow$ Fleurs $7.20/3.94$ N/A N/A *$2.62/2.69$ N/A $\mathbf{13.4/4.56}$
AST (Ko2En) BLEU/Score (1-5) $\uparrow$ Fleurs $14.0/3.72$ $18.6/3.25$ N/A *$10.4/2.80$ N/A $\mathbf{19.2/4.67}$
AST (Ko2En) BLEU/Score (1-5) $\uparrow$ Minds14 $26.3/4.12$ $\mathbf{29.5}/4.00$ N/A *$14.8/3.18$ N/A $26.3/\mathbf{4.52}$
SSUM (En) Score (1-7) $\uparrow$ SDS-PART6 $5.40$ N/A $4.54$ $5.30$ $5.48$ $\mathbf{5.87}$
SSUM (Ko) Score (1-7) $\uparrow$ KMSS $4.12$ N/A N/A N/A N/A $\mathbf{4.37}$
SQQA (En) ACC $\uparrow$ LibriSQA $\mathbf{69.5}$ N/A $57.2$ $64.5$ $48.9$ $67.0$
SQQA (Ko) ACC $\uparrow$ KorQuAD- speech $\mathbf{64.9}$ N/A N/A N/A N/A $64.8$
@@
  • ASR: Achieves a SOTA CER of 3.82 on Korean benchmarks and a competitive WER of 6.83 on English OpenASR.
  • AST: Showcases a preference for semantic quality over n-gram overlap, scoring higher in GPT-4 evaluation than Whisper on the Minds14 Ko2En task despite a lower BLEU score.
  • SSUM: Delivers competitive scores of 5.40 (English) and 4.12 (Korean), performing on par with several larger models.
  • SQQA: Sets the SOTA on all tested benchmarks, with leading accuracies of 69.5% on LibriSQA (English) and 64.9% on KorQuAD-speech (Korean).
In summary, FastSLM provides a highly effective and efficient solution for diverse speech-language tasks, proving that a compact speech representation can drive s SOTA performance.
@@ 4.5 Ablation Study
Effect of Hierarchical Staging in HFQ-Former: To directly evaluate the benefit of the hierarchical design in HFQ-Former, we compare three variants using only Stage 1, Stage 1–2, and the full Stage 1–2–3. Table 4 shows that performance consistently improves as more hierarchical stages are included, demonstrating that progressive temporal abstraction is essential for long-form speech processing.
Stage 1 2.23 4.12 56.7
Stage 1/2 2.15 4.98 62.2
Stage 1/2/3 2.09 5.40 64.9
@@
Our analysis identifies 1.67 tokens/sec as an optimal operating point that balances these competing factors, achieving strong performance while minimizing computational demands. This decision is supported by a mathematical analysis of the point of diminishing returns, detailed in Appendix I.
Scaling Limits of Long-Form Speech Input with FastSLM: To provide a practical assessment beyond indirect complexity metrics like parameters and FLOPs (Ma et al., 2018), we empirically evaluate the scaling properties of FastSLM. We measure two key indicators on a single 40GB NVIDIA A100 GPU: VRAM consumption to assess memory efficiency and time-to-first-token (TTFT) to quantify the latency introduced by our multi-layered HFQ-Former. The results, presented in Fig. 4, highlight significant advantages in scalability. While benchmark models exhibit exponential VRAM growth, FastSLM demonstrates near-linear scaling, successfully processing an 8 hours (28,800 seconds) speech stream using under 30GB of memory. Regarding latency, FastSLM maintains a competitive TTFT. Although its initial latency is comparable to the similarly-sized Voxtral-Mini, it scales far more effectively, showing only a minimal increase as the speech length grows, in contrast to the sharp rise observed in the baseline. These findings confirm that FastSLM architecture enables robust inference on ultra-long-form speech far beyond the capabilities of existing models, especially within constrained environments. Furthermore, its efficiency is highly advantageous for batch processing; the minimal memory footprint per stream allows for significantly larger batch sizes on a single GPU, thereby maximizing throughput for parallelized workloads.
@@
Effect of Hierarchical Modules and Training Stage: To validate the effectiveness of our architecture and three-stage training strategy, we performed an ablation study by selectively removing key components: the Downsampler stage, the hierarchical attention mechanism, and Training Stage 2 (dedicated to long-form speech adaptation). The results, summarized in Table 5, highlight the critical contribution of each component. Removing the Downsampler stage or hierarchical attention leads to substantial performance degradation across all benchmarks—WER on LS-Long (Park et al., 2024) rises to 12.4 and 10.8, respectively. More importantly, omitting Training Stage 2, which enhances long-context understanding through ASR-based pretraining, results in notable drops across tasks, increasing LS-Long WER from 5.98 to 6.81 and reducing KorQuAD-Speech accuracy from 64.9% to 59.0%. The full FastSLM configuration consistently outperforms all ablated variants, demonstrating that both the hierarchical architectural design and the dedicated Training Stage 2 are indispensable for achieving efficient and robust long-form speech adaptation.
@@
FastSLM $\mathbf{5.98}$ $\mathbf{64.9}$ $\mathbf{5.40}$
@@ 5 Conclusion
In this paper, we introduce FastSLM, a lightweight and efficient SLM designed to overcome the critical scaling limitations of processing long-form speech. At the core of our approach is the HFQ-Former, a novel module that hierarchically compresses high-frame speech features into an optimal representation while preserving both local and global context. This architecture is complemented by a cost-effective three-stage training strategy, which enables robust adaptation of a pre-trained LLM for the speech modality. FastSLM achieves SOTA or competitive performance across diverse benchmarks while reducing the high-frame level features of conventional speech encoders by up to 97%, dramatically lowering GPU memory usage and computational cost without sacrificing accuracy. These results demonstrate that FastSLM provides a practical and scalable solution for efficient long-form speech understanding. By enabling models to process and reason over one of the primary modalities of human communication, this work contributes a critical building block for future multimodal systems aspiring toward AGI. Despite these promising results, we acknowledge several limitations of our current approach. A detailed discussion of these, along with potential directions for future research, is provided in Appendix J.
@@ Appendix B Additional Analysis of Hierarchical Temporal Modeling
Appendix C Prompt Template and Hierarchical Tags for Training
@@
Appendix D Pre-training Dataset
@@
Total 15,018 10,212,348 -
Appendix E Prompt for GPT-4 as a Judge on Speech Benchmarks
The following is the exact prompt template used for evaluating AST output via an LLM-as-a-Judge. Placeholders like SOURCE_TRANSCRIPT are filled in programmatically for each evaluation instance.
Listing 1: LLM-as-a-Judge Prompt for AST Evaluation
# [Role]
You are a highly skilled professional evaluator specializing in {speech_language} to{target_language} Automatic Speech Translation (AST). Your task is to meticulously evaluate the quality of a {target_lang} translation generated from an English speech transcript.
Critically, you must recognize that the source text is not a manually written sentence but a transcript generated by an Automatic Speech Recognition (ASR) system. Therefore, the source itself may contain errors.
# [Evaluation Criteria]
You will conduct your evaluation based on the following two core criteria and an error analysis.
# 1. Adequacy: Accuracy of Meaning
Rate on a scale of 1-5 how accurately the core meaning and nuances of the source text are preserved in the target translation.
- **5 (Excellent):** All information and nuances from the source are perfectly conveyed without any loss or distortion.
- **4 (Good):** The core meaning is fully conveyed, but some minor nuances or details are lost.
- **3 (Fair):** The core meaning is conveyed, but some important information is missing or translated slightly inaccurately.
- **2 (Poor):** The subject, object, or key terms are mistranslated, leading to a significant distortion of the sources meaning.
- **1 (Very Poor):** The translation is a complete mistranslation and does not convey any of the sources meaning.
# 2. Fluency: Naturalness of the Translation
Rate on a scale of 1-5 how grammatically correct and natural the target translation sounds to a native Korean speaker.
- **5 (Excellent):** The translation is grammatically perfect and sounds completely natural, as if written by a native speaker.
- **4 (Good):** The translation is grammatically correct but feels slightly unnatural or "like a translation."
- **3 (Fair):** The meaning is understandable, but there are clear grammatical errors or awkward expressions.
- **2 (Poor):** The sentence structure is very awkward or contains many grammatical errors, making it difficult to understand.
- **1 (Very Poor):** The output is a collection of words that is not comprehensible as a sentence.
# [Critical Instructions for AST Evaluation]
- **Potential for ASR Errors:** If a word in the source transcript seems highly out of place in its context (e.g., the word bricks in a description of a climbing destination), it is likely an ASR error.
- **Evaluating Error Correction:** If the translation model **corrects** such a likely ASR error into a contextually appropriate term (e.g., translating bricks as the Korean term for granite walls, this is a highly positive attribute. You must mention this in the adequacy_reason and award a high score for Adequacy.
# [Inputs]
- **ASR Transcript (Source):** {SOURCE_TRANSCRIPT}
- **Model Translation (Candidate):** {MODEL_TRANSLATION}
- **Human Reference (Gold):** {HUMAN_REFERENCE}
# [Output Format]
You MUST provide your evaluation results in the following JSON format.
adequacy_score: <Float, 1.0-5.0>,
fluency_score: <Float, 1.0-5.0>,
overall_score: <Float, 1.0-5.0>
Listing 2: LLM-as-a-Judge Prompt for SSUM Evaluation
@@ Appendix E Prompt for GPT-4 as a Judge on Speech Benchmarks
explanation: "Brief reasoning for the score based on the rubric.",
@@
Appendix F Generation and Decoding Configuration

### v2

@@ 4.1 Dataset Description
Pre-training Dataset: As described in Section 3.2, we constructed a bilingual dataset comprising 17K hours speech-text pairs to adapt ASR capabilities to the LLM during the pre-training stage. Including 9,152 hours of English speech-text pair (LibriSpeech (Panayotov et al., 2015), GigaSpeech-L (Chen et al., 2021), Voxpopuli (Wang et al., 2021), SpgiSpeech-M (O’Neill et al., 2021), Earnings-22 (Rio et al., 2022), AMI (Kraaij et al., 2005), Common Voice 15 (Ardila et al., 2019), AI-HUB ASR-En (The Open AI Dataset Project, 2021)) dataset, and 7,812 hours of Korean speech-text pair (AI-Hub-ASR-Ko (The Open AI Dataset Project, 2021)) dataset. A detailed description of the pre-training dataset can be found in Appendix E Table 7.
Long-form speech Dataset: To enhance the model capacity to process long-form speech input, particularly for tasks such as SSUM and SQQA, we constructed a dedicated long-form speech dataset. For English, we curated a total of 1,224 hours of long-form speech. For Korean, we curated a total of 1,012 hours of long-form speech.
@@
  • AST: We evaluated translation on the Fleurs (En2Ko/Ko2En) (Conneau et al., 2022) and Minds14 (Ko2En) (Gerz et al., 2021) datasets. We measured translation performance using the BLEU score (Post, 2018).
  • SSUM: Evaluation was conducted on SDS-PART6 (He et al., 2025) and KMSS-speech (Kim et al., 2022). Summarization quality was assessed using GPT-4 scoring with the LLM-as-a-judge framework (Zheng et al., 2023). Please refer to Appendix F for the SSUM judge prompt for GPT-4.
@@ 4.2 Experimental Setup
Model Architecture: FastSLM employs the encoder from Whisper-large-v3 (Radford et al., 2023) for speech feature extraction and adopts Qwen3-4B (Yang et al., 2025) as the backbone LLM for text generation. Despite its relatively compact size, Qwen3-4B exhibits sufficient capacity for comprehending speech-derived representations. In contrast to prior SLMs that typically utilize LLM backbone with 7 to 14 billion (B) parameters (Chu et al., 2023, 2024; Tang et al., 2024; Yu et al., 2024; Rubenstein et al., 2023; Ding et al., 2025; Liu et al., 2025a), FastSLM achieves a favorable cost-performance trade-off by leveraging lightweight architecture without compromising performance (Abouelenin et al., 2025; Ghosh et al., 2025). The HFQ-Former module within FastSLM compresses frame-level features via a hierarchical query-based mechanism. The number of learnable queries of $\mathbf{Q}_{i}^{A}$ was set to 80 cost-effectively (Yu et al., 2024), and the number of learnable queries of $\mathbf{Q}_{c}^{A}$ used as contextual input in LLM was set to 50 through the experiment. For further ablation studies and justifications for parameter selection, please refer to Section 4.5 Fig. 3.
@@
The model implementation details and the training setup are summarized in Appendix H.
@@ 4.3 Comparison with Baseline Q-Former
HFQ-Former (ours) $2.09$ $4.67$ $\mathbf{6.55}$ $\mathbf{1.67}$ $\mathbf{2.15}(-92\%)$
@@
In contrast, HFQ-Former achieves a strong balance between accuracy and efficiency: it reduces the token rate by 38% compared to WQ-Former and lowers LLM FLOPs by 41.1% (3.65T → 2.15T), while still improving WER on VoxPopuli. These results indicate that HFQ-Former provides a significantly more efficient speech-to-LLM alignment mechanism, greatly reducing the computational burden of autoregressive decoding for long-form speech without compromising recognition quality.
Table 3: Comparison of FastSLM with other SLMs on various tasks. Bold indicates the best performance, and underline indicates the second-best performance. WER and CER are lower-is-better ($\downarrow$); Accuracy (ACC), BLEU, and Score are higher-is-better ($\uparrow$). N/A indicates the model lacks the corresponding capability. ‘*’ denotes results fine-tuned on an additional Korean dataset provided in the official Hugging Face supplementary material. Detailed ASR benchmark WER for each dataset is reported in Appendix I.
#speech tokens/30 Sec. 50 1500 101 375 750 960
ASR (En) WER $\downarrow$ OpenASR $\underline{6.47}$ $7.44$ $7.43$ $\mathbf{6.14}$ $7.05$ $9.29$
ASR (Ko) CER $\downarrow$ Fleurs Common Voice 15 $\mathbf{3.82}$ $7.92$ N/A N/A N/A $\underline{4.55}$
AST (En2Ko) BLEU $\uparrow$ Fleurs $\underline{7.39}$ N/A N/A *$2.62$ N/A $\mathbf{13.4}$
AST (Ko2En) BLEU $\uparrow$ Fleurs $\mathbf{19.5}$ $18.6$ N/A *$10.4$ N/A $\underline{19.2}$
AST (Ko2En) BLEU $\uparrow$ Minds14 $\underline{28.9}$ $\mathbf{29.5}$ N/A *$14.8$ N/A $26.3$
SSUM (En) Score (1-7) $\uparrow$ SDS-PART6 $5.40$ N/A $4.54$ $5.30$ $\underline{5.48}$ $\mathbf{5.87}$
SSUM (Ko) Score (1-7) $\uparrow$ KMSS $\underline{4.12}$ N/A N/A N/A N/A $\mathbf{4.37}$
SQQA (En) ACC $\uparrow$ LibriSQA $\mathbf{69.5}$ N/A $57.2$ $64.5$ $48.9$ $\underline{67.0}$
SQQA (Ko) ACC $\uparrow$ KorQuAD- speech $\mathbf{64.9}$ N/A N/A N/A N/A $\underline{64.8}$
@@ 4.4 Quantitative Results
  • ASR: Achieves a SOTA CER of 3.82 on Korean benchmarks and a competitive WER of 6.47 on English OpenASR.
  • AST: Achieves SOTA performance on the Fleurs Ko2En task with 19.5 BLEU, outperforming both Whisper and Gemini-2.5-Flash. Furthermore, it demonstrates competitive performance on the Minds14 dataset (28.9 BLEU), comparable to the strong Whisper baseline.
  • SSUM: Delivers competitive scores of 5.40 (English) and 4.12 (Korean), performing on par with several larger SLMs.
  • SQQA: Achieves SOTA performance on the evaluated benchmarks, attaining accuracies of 69.5% on LibriSQA (English) and 64.9% on KorQuAD-speech (Korean)
In summary, FastSLM provides a highly effective and efficient solution for diverse speech-language tasks, proving that a compact speech representation can drive SOTA performance.
@@ 4.5 Ablation Study
Our analysis identifies 1.67 tokens/sec as an optimal operating point that balances these competing factors, achieving strong performance while minimizing computational demands. This decision is supported by a mathematical analysis of the point of diminishing returns, detailed in Appendix J.
@@
Scaling Limits of Long-Form Speech Input with FastSLM: To provide a practical assessment beyond indirect complexity metrics like parameters and FLOPs (Ma et al., 2018), we empirically evaluate the scaling properties of FastSLM. We measure two key indicators on a single 40GB NVIDIA A100 GPU: VRAM consumption to assess memory efficiency and time-to-first-token (TTFT) to quantify the latency introduced by our multi-layered HFQ-Former.
The results, presented in Fig. 4, highlight significant advantages in scalability. While benchmark models exhibit exponential VRAM growth, FastSLM demonstrates near-linear scaling, successfully processing an 8 hours (28,800 seconds) speech stream using under 30GB of memory. Regarding latency, FastSLM maintains a competitive TTFT. Although its initial latency is comparable to the similarly-sized Voxtral-Mini, it scales far more effectively, showing only a minimal increase as the speech length grows, in contrast to the sharp rise observed in the baseline. These findings confirm that FastSLM architecture enables robust inference on ultra-long-form speech far beyond the capabilities of existing models, especially within constrained environments. Furthermore, its efficiency is highly advantageous for batch processing; the minimal memory footprint per stream allows for significantly larger batch sizes on a single GPU, thereby maximizing throughput for parallelized workloads.
Impact of Hierarchical Modules and Training Strategy: To validate the efficacy of our architectural design and the three-stage training strategy, we conducted an ablation study by selectively removing key components: the Downsampler stage, the hierarchical attention mechanism, and Training Stage 2 (long-form speech adaptation). As summarized in Table 4, removing the Downsampler or hierarchical attention caused a severe performance degradation, with WER on LS-Long (Park et al., 2024) surging to 12.4 and 10.8, respectively. This underscores the structural necessity of our hierarchical compression in handling long sequences. Furthermore, omitting Training Stage 2, which aligns the model with long-context speech via Long-form speech adaptation, resulted in significant regression across tasks—specifically, increasing LS-Long WER from 5.98 to 6.81 and dropping KorQuAD-Speech accuracy from 64.9% to 59.0%. The full FastSLM configuration consistently outperforms all ablated variants, confirming that both the hierarchical architecture and the dedicated adaptation stage are integral to achieving efficient and robust long-form understanding.
@@
FastSLM $\mathbf{5.78}$ $\mathbf{64.9}$ $\mathbf{5.40}$
@@ 5 Conclusion
In this paper, we introduced FastSLM, a lightweight and efficient SLM engineered to overcome the critical scaling limitations of long-form speech processing. At the core of our approach is the HFQ-Former, a novel module that hierarchically compresses high-frame-rate speech features into compact semantic tokens. Complemented by a cost-effective three-stage training strategy, FastSLM achieves SOTA or competitive performance across diverse benchmarks while reducing feature sequence length by up to 96%. These results demonstrate that extreme token compression is a practical and scalable solution for real-time long-form speech understanding, contributing a critical building block for future multimodal systems. A discussion on current limitations and future directions is provided in Appendix K.
@@ Appendix B Additional Analysis of Hierarchical Temporal Modeling
Appendix C Effect of Hierarchical Staging in HFQ-Former
To directly evaluate the benefit of the hierarchical design in HFQ-Former, we compare three variants using only Stage 1, Stage 1–2, and the full Stage 1–2–3. Table 6 shows that performance consistently improves as more hierarchical stages are included, demonstrating that progressive temporal abstraction is essential for long-form speech processing.
Stage 1 2.23 4.12 56.7
Stage 1/2 2.15 4.98 62.2
Stage 1/2/3 2.09 5.40 64.9
Appendix D Prompt Template and Hierarchical Tags for Training
@@
Appendix E Pre-training Dataset
@@
Total 16,964 10,212,348 -
Appendix F Prompt for GPT-4 as a Judge on Speech Benchmarks
The following is the exact prompt template used for evaluating SSUM output via an LLM-as-a-Judge.
Listing 1: LLM-as-a-Judge Prompt for SSUM Evaluation
@@ Appendix F Prompt for GPT-4 as a Judge on Speech Benchmarks
explanation: Brief reasoning for the score based on the rubric.,
@@
Appendix G Generation and Decoding Configuration